OpenStudy (anonymous):

simplify: 5√(-32y^5)^2

OpenStudy (saifoo.khan):

\[5\sqrt{(-32y^5)^2 }\] like that?

OpenStudy (anonymous):

I think so, the first 5 is ^5

OpenStudy (anonymous):

I think, I'mnot familiar with these...

OpenStudy (anonymous):

No wait, that does seem correct :)

OpenStudy (saifoo.khan):

\[\sqrt[5]{(-32y^5)^2 }\]

OpenStudy (saifoo.khan):

like that^

hero (hero):

Swap the 2 with the 5: \[\sqrt[5]{(-2^2y^2)^5}\]

hero (hero):

Then cancel the root and the 5th power

hero (hero):

\[4y^2\] should be the answer

OpenStudy (anonymous):

Thanks Hero... but apparently it's not the answer we are looking for.